Database last updated on Sun, 01 Feb 2026 17:40:48 +0000 by localhost

Instrument constant changes for the week before 2023253 (2 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 23 rows)

Measurements and tests summary.


_jday_loc_serial_skc
253TORONTO#026menu
(0 sec)

Dead time (0 sec)

Nothing to report.

Standard lamp R6 (1 sec)

Nothing to report.

Standard lamp R5 (1 sec)

Nothing to report.

Number of days since last dome/window cleaning (0 sec, 2 rows)

serialdays
026more than a year

Missing files (0 sec, 1 rows)

Seriallast datadays ago
02620232540-1

Run/Stop results for last week. (0 sec, 1 rows)

SerialTotalFailed 0Failed darkFailed 1Failed 2Failed 3Failed 4Failed 5Failed double
02611300042011

Ozone results for good DS obs (std<2.5DU, mu<4) within 3 minutes between 2 Brewers. (1 sec)

No ozone measurements to display.

Unsorted comments (0 sec, 10715 rows)

SerialtypeTextCountYYYYJJJ
026env:DosBox12023253
026cubdsp:012023253
026fre_after:da_lo 9633 963992023253
026fre_after:da_lo 9652 9658152023253

The END (0 sec, 2 rows)

If you have questions and/or comments about this page please contact Volodya Savastiouk